OpenStudy (anonymous):

If cosx = 3/4 and 0< x < pi/2. Find sin x/2

OpenStudy (anonymous):

cosx=1-2sin^2(x/2) hence sinx/2=sqrt(1-cosx)/2 =sqrt(1/8)

OpenStudy (anonymous):

Thanks!